Circular No. 7576

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001P IN NGC 3947 W. D. Li, University of California at Berkeley, on behalf of LOTOSS

(cf. IAUC 7514), reports the discovery of a supernova (mag ∼ 17.5) on unfiltered images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Jan. 31.4 and Feb. 1.3 UT. SN 2001P is located at α = 11^{h}53^{m}21^{s}.56, δ = +20^{o}44′53′′.3 (equinox 2000.0), which is 17′′.2 east and 13′′.1 south of the nucleus of NGC 3947. A KAIT image taken on Jan. 7.4 showed nothing at this position (limiting mag ∼ 19.0).

A. V. Filippenko, R. Chornock, and M. Modjaz, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Feb. 1 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2001P is a type-Ia supernova, roughly 1 week past maximum brightness. Some spectral features (e.g., Ti II absorption near 430 nm) suggest that it might be subluminous.

SUPERNOVA 2001Q IN UGC 6429 Li also reports the discovery of another supernova on unfiltered KAIT

images taken on Jan. 31.4 (mag ∼ 18.6) and Feb. 1.3 UT (mag ∼ 18.3). SN 2001Q is located at α = 11^{h}25^{m}19^{s}.77, δ = +63^{o}43′15′′.6 (equinox 2000.0), which is 4′′.6 east and 29′′.6 south of the nucleus of UGC 6429. A KAIT image of the same field taken on Jan. 2.4 showed nothing at this position (limiting mag ∼ 19.0).

Filippenko et al. note that spectra obtained as above (but with poor quality longward of 520 nm) indicate that SN 2001Q is a young type-II supernova; the hydrogen Balmer absorption lines are prominent.

SUPERNOVA 2001M IN NGC 3240 N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), writes:

“Spectra (range 390–680 nm; resolution 1000) of SN 2001M (cf. IAUC 7568) were taken (by J. Colianni, T. Gokas, and F. Winkler, Middlebury College; and R. C. Smith, CTIO) with the CTIO 1.5-m telescope facility spectrograph on Jan. 25.3 and 26.3 UT. The reduced spectrophotometry shows that SN 2001M is a type-Ic supernova similar to SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11), roughly 1–2 weeks past maximum light. Si II at 615 nm is weak or absent, and there is no evidence of prominent He I lines. The Na I absorption at 589 nm is also weak, indicating that it is not more than a few weeks past maximum light.”

2001 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green